Two incinerators are being considered by a waste management company. Design A has an initial cost of $2,825,000, has annual operating and maintenance costs of $950,000, and requires overhauls every 5 years at a cost of $1,100,000. Design B is more sophisticated, including computer controls; it has an initial cost of $5,400,000, has annual operating and maintenance costs of $400,000, and requires overhauls every 10 years at a cost of $2,800,000.

Using a 8.0 %/year interest rate, determine the capitalized cost for each design.

Design A: $ ?
Design B: $ ?

Recommend which design should be chosen:
Answer #1

Design A=2825000+950000/8%+1100000/1.08^5*1/(1-1/1.08^5)=17043776.25

Design B=5400000+400000/8%+2800000/1.08^10*1/(1-1/1.08^10)=12816032.10

Design B should be chosen as it has lower capitalized cost